in the following models, paying attention to the functional form, how would you interpret the coefficient b1;

(I) log(price)=b0+b1log(nox) +b2 rooms +U

(ii) In(y)= b0 +b1 x +u

(iii) y=b0 +b1 In(x) +u

(iii) In(y)=b0 +b1 In(x) + u
Answer #1

Answer

(I)

log(price)=b0+b1log(nox) +b2 rooms +U

b1 will be interpreted as follows :

1% increase in nox will result in increase in Price by b1%

(ii)

In(y)= b0 +b1x +u

b1 will be interpreted as follows

1 unit increase in x will result in increase in y by (100*b1)%

(iii)

y=b0 +b1 In(x) +u

b1 will be interpreted as follows

1% increase in x will result in increase in y by (b1/100) units.

(iv)

In(y)=b0 +b1 In(x) + u

b1 will be interpreted as follows

1% increase in x will result in increase in y by b1%.